OpenStudy (anonymous):

What is the slope of the line 4y – 6x + 12 = 0 in simplest form? will fan and medal

OpenStudy (anonymous):

4y - 6x + 12 We can rearrange into slope-intercept form. 4y - 6x + 12 Subtract 12 to both sides: 4y - 6x = -12 Add 6x to both sides, what do you get?

OpenStudy (anonymous):

4y=-6

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

We are not combining anything because there are no like terms, try again.

OpenStudy (anonymous):

4y= 6x-12?

OpenStudy (anonymous):

Yes! Now divide 4 to both sides.

OpenStudy (anonymous):

4y divided by 12 would be 3 and that would be 6x-3?

OpenStudy (anonymous):

Yes, but we also divide it to 6.

OpenStudy (anonymous):

That gives us 6/4x

OpenStudy (anonymous):

SO we have: y = 6/4x - 3 Can you simplify 6/4? Divide the numerator and denominator by 2.

OpenStudy (anonymous):

*So

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

just remember y=mx+b m is the slope

OpenStudy (anonymous):

Yes, so the slope is 3/2.

OpenStudy (holly00d1248):

I agree.